# Privacy Risks and Preservation Methods in Explainable Artificial Intelligence: A Scoping Review

### v2

@@ 2.3.4 Feature-based methods
These explanation methods score or measure the effect of individual input features on the output of the model (Arrieta et al., 2020; Bhatt et al., 2020; Dwivedi et al., 2023; Strobel & Shokri, 2022). They are also referred to as feature importance (McDermid et al., 2021), feature relevance (Arrieta et al., 2020) or attribution-based (Liu et al., 2024) methods. They are based on the attribution problem which is the distribution of the output of a model for a specific input to its base features (Sundararajan & Najmi, 2020). Two important categories of feature-based methods identified in literature are perturbation and backpropagation-based methods (Ancona et al., 2018; McDermid et al., 2021).
  • Perturbation-based methods remove, alter, or mask an input feature or set of features and observe the difference with the original output (McDermid et al., 2021). Some perturbation-based methods are LIME (Ribeiro et al., 2016), permutation feature importance (Breiman, 2001), SHAP (Lundberg & Lee, 2017) and MASK (Fong & Vedaldi, 2017).
  • Backpropagation-based methods compute input attributions in forward and backward passes of the network (Ancona et al., 2018). The use of the gradient of the output with the respective input features (McDermid et al., 2021; Strobel & Shokri, 2022) is a common approach in these methods and is referred to as gradient-based approach. Methods used on images that determine the global importance of pixels, generate saliency maps, and are referred to as pixel-level attribution methods (Kapishnikov et al., 2019; Molnar, 2023). Some examples of backpropagation-based methods are gradient (Simonyan et al., 2014), gradient x input (Shrikumar et al., 2017b), guided backpropagation (Springenberg et al., 2015) and integrated gradients (Sundararajan et al., 2017).
@@ 2.3.5 Concept-based methods
Concept-based methods aim to uncover high-level concepts above features, pixels and characters (Ghorbani et al., 2019), that influence the behaviour of a model (Bereska & Gavves, 2024). The use of human understandable concepts result in naturally interpretable explanations. Desiderata such as meaningfulness, coherence and importance are proposed for this category of explanations (Ghorbani et al., 2019). Concept activation vectors (CAVs) (Kim et al., 2018) is a concept-based method that leverages linear classifiers to determine the presence or absence of concepts corresponding to a set of input examples. The use of concept activation regions (Crabbé & van der Schaar, 2022) enhances CAVs and allows the underlying examples to be distributed across the latent space of the model, resulting in the discovery of global explanations. In another approach, concept bottleneck models (Koh et al., 2020) predict intermediate human-specified concepts in the model and utilise these for making predictions. The models also support human interventions on these concepts and facilitate the generation of counterfactuals, thus enabling human-model collaboration (Koh et al., 2020).
@@ 2.3.6 Probing-based methods
Probing-based methods use classifiers or probes to determine the knowledge captured by a model (Sajjad et al., 2022; Schneider, 2024a). The technique involves using the internal activations of networks to train probing classifiers for prediction of embedded properties (Alain & Bengio, 2017; Belinkov, 2022). However, such probing experiments require the properties to be known apriori along with clear specifications of the original/probing tasks, the models and datasets used. The outcomes of these experiments can be applied for tuning the model on certain tasks or determining the encoded information for downstream tasks.
@@ 2.3.7 Neuron activation methods
These methods focus on the behaviour of neurons that are responsible for specific outcomes or represent learned linguistic properties (Zhao et al., 2024). Proposed methods in this category, such as linguistic correlation analysis and cross-model correlation analysis, identify and analyse such neurons for post-hoc explainability (Dalvi et al., 2019). Individual or groups of neurons may be discovered as salient for a property and ranked based their importance to the model’s task (Dalvi et al., 2019). These methods may also experience trade-offs between accuracy and selectivity (Antverg & Belinkov, 2022).
@@ 2.3.8 Mechanistic interpretability
Mechanistic interpretability (Bereska & Gavves, 2024) seeks to understand the model outputs by reverse engineering (Olah, 2022) the system. Analogous to understanding complex computer programs, this approach aims at dissecting network activations into independently understandable units. It considers the analysis of individual components of the system, such as features, connections and neurons, to determine the causes of outputs. This approach to interpretability includes techniques that could be applied before, during or after the training process. Techniques, such as feature visualisation (Zimmermann et al., 2021) and sparse autoencoders (Cunningham et al., 2024), are some of the methods from this category.
@@
Example-based Anchors Model-agnostic Ribeiro et al. (2018)
Contrastive explanations Model-agnostic Dhurandhar et al. (2018)
Counterfactuals Model-agnostic Wachter et al. (2017)
Influence functions Model-agnostic Koh & Liang (2017)
Prototypes and criticisms Model-agnostic Kim et al. (2016)
Knowledge-based Semantic web technologies Model-agnostic Seeliger et al. (2019)
Neuro-symbolic approaches Model-specific Hitzler et al. (2022)
Feature-based Perturbation-based LIME Model-agnostic Ribeiro et al. (2016)
Permutation Feature Importance Model-agnostic Breiman (2001)
SHAP Model-agnostic Lundberg & Lee (2017)
MASK Model-agnostic Fong & Vedaldi (2017)
Backpropagation-based Gradient Model-specific Simonyan et al. (2014)
Gradient x Input Model-specific Shrikumar et al. (2017b)
Guided Backpropagation Model-specific Springenberg et al. (2015)
Integrated Gradients Model-specific Sundararajan et al. (2017)
Concept-based Concept activation vectors Model-specific Kim et al. (2018)
Concept bottleneck models Model-specific Koh et al. (2020)
Concept activation regions Model-specific Crabbé & van der Schaar (2022)
Probing-based Probing experiments Model-specific Alain & Bengio (2017)
Neuron activation Linguistic correlation analysis Model-specific Dalvi et al. (2019)
Cross-model correlation analysis Model-specific Dalvi et al. (2019)
Mechanistic interpretability Feature visualisation Model-specific Zimmermann et al. (2021)
Sparse autoencoders Model-specific Cunningham et al. (2024)
@@ 2.4 Related reviews
An in-depth overview of the core concepts and taxonomies in XAI was provided by Arrieta et al. (2020). Mohseni et al. (2021) conducted an interdisciplinary survey and proposed a comprehensive framework for design and evaluation of XAI methods. Dwivedi et al. (2023) covered a wide breadth of explanation algorithms, programming frameworks and software toolkits for XAI development. Ali et al. (2023) examined explainability through the lens of trustworthiness detailing evaluation metrics, available software packages and XAI datasets. Bodria et al. (2023) systematically categorized explanation methods and benchmarked prominent methods using quantitative metrics. Muralidhar et al. (2023) reviewed transparency elements from human computer interaction (HCI) in the context of explanations while Cambria et al. (2023) investigated presentation methods and usage of natural language with XAI.
@@
Beyond these broad surveys, domain specific reviews have also emerged. For example, XAI in healthcare has been surveyed by Payrovnaziri et al. (2020) and Yang et al. (2022); in cybersecurity by Capuano et al. (2022) and in energy and power systems by Machlev et al. (2022). Methodology focussed reviews also exist, covering counterfactuals (Guidotti, 2022), data-driven knowledge-aware XAI systems (Li et al., 2020b), knowledge-graph based XAI (Rajabi & Etminani, 2022; Tiddi & Schlobach, 2022) and semantic web technologies for explanations (Seeliger et al., 2019). Recent advances include the intersection of explainability and federated learning (FL), termed as Federated XAI (Fed-XAI), reviewed by López-Blanco et al. (2023) and categorisation of explanation techniques for transformer-based language models based on training paradigms as surveyed by Zhao et al. (2024). A comprehensive review of mechanistic interpretabililty in LLMs for achieving AI safety was conducted by Bereska & Gavves (2024).
@@
Recent literature has increasingly highlighted the potential of malicious exploitation of XAI interfaces. Baniecki & Biecek (2024) present a survey of adversarial attacks in XAI and discuss corresponding defenses. While their study addresses a specific type of XAI privacy attack, its primary emphasis is on non-privacy attacks including data poisoning, backdoor attacks, model manipulation and attacks on fairness metrics through explanations. A categorisation of XAI-aware attacks was proposed in another recent work (Noppel & Wressnegger, 2024b; a) based on the outputs that are preserved in the attack. The work also categorised attacks based on the scope of explanation alteration and adversary capabilities. However, this work targets the robustness of post-hoc XAI and though the proposed categorisation is relevant to privacy attacks, it does not directly focus on the privacy aspect of XAI safety.
@@
The focus of this review diverges from prior reviews by specifically examining the privacy risks that target the identification and exposure of personal or sensitive information through the misuse of XAI interfaces in AI systems. Further, we review strategies used by researchers in mitigating the privacy leakage in XAI. An early work by Spartalis et al. (2023), provides a short review on XAI privacy risks and applicable defense mechanisms. The review also identifies security risks such as evasion and poisoning through XAI interfaces. Though this work raises awareness for further research in this area, it mainly focuses on early privacy attacks and discusses limited defense mechanisms. A recent review (Nguyen et al., 2025) explores the intersection of privacy and explainability through attacks and countermeasures, but differs from our review in its methodology and lacks a detailed examination of the core requirements underpinning the design of privacy preserving XAI. Our review employs an established scoping review methodology guided by clearly defined research questions. The resulting taxonomy of XAI privacy risks and corresponding mitigation methods are distilled from the understanding of existing literature across the privacy and XAI communities. This methodology enables a structured and rigorous approach to addressing the research questions through the analysis of the selected studies.
@@ 3 Method
We conducted a scoping review based on the Preferred Reporting Items for Systematic reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) (Tricco et al., 2018). This section elaborates the process followed, the proposed taxonomy of XAI privacy risks and the identified research trends.
@@ 3.1 Literature selection and extraction
A 4-step process was employed comprising of identification, screening, eligibility, and extraction, as illustrated in Figure 2. In the initial step of identification, Elsevier Engineering Village (Engineering Village, ) search platform was used and the search was conducted on Compendex and Inspec databases. These databases index publications from leading computer science publishers, including IEEE, ACM, Springer and Elsevier. A researcher formulated the search string, using the two main concepts of privacy and explainability, with the help of a librarian. This was applied on the title, subject, and abstract fields. The reseacher worked with the librarian to set the search, inclusion and exclusion criteria as detailed below for reproducibility:
@@
  • Period of publication: January 1, 2019, to December 31, 2024. The start year was chosen based on the seminal works (Milli et al., 2019; Shokri et al., 2020; 2021) published on this topic.
@@
The search results comprising of 3,766 studies were exported from Engineering Village and imported into Covidence (Covidence, ) review management software by one researcher. During the import process, the software merged duplicate studies from the databases, retaining only unique records. After deduplication, 1,943 studies were forwarded for screening wherein the title and abstract were examined by the researcher to determine relevance to RQ1 or RQ2 while considering the inclusion criteria. Out of 1,943 studies, 69 studies were moved to the next step to determine eligibility wherein the full text of the identified articles were examined with respect to RQ1 and RQ2. This step was conducted by one researcher and independently verified by a second researcher. Since the focus of this article is privacy leakage through XAI interfaces, studies that addressed adversarial XAI (such as backdoor, poisoning, etc.) without overlap on privacy were eliminated. Studies that discussed general privacy issues in ML without XAI context were also eliminated. Related survey papers that appeared in the search results were not extracted but used to determine any additional studies relevant to the research questions. These steps resulted in removal of 17 studies and addition of 5 studies identified through forward and backward searches. The removal and addition of studies were discussed and verified by two researchers. Thus, overall the 4-step scoping review process resulted in extraction of 57 studies.
@@ 4 Privacy Risks in XAI
Traditionally privacy is referred to as the “right to be left alone” (Warren & Brandeis, 1890) and the “claim of individuals, groups, or institutions to determine for themselves when, how, and to what extent information about them is communicated to others” (Westin, 1967). In the modern context, with availability, collection, and collation of copious information about individuals through online and offline sources, the concept of information privacy is more applicable and refers to the ability of individuals to exert control on their own data (Curzon et al., 2021). Clarke (1999) has defined information privacy as the “claims of individuals that data about themselves should generally not be available to other individuals and organizations, and that, where data is possessed by another party, the individual must be able to exercise a substantial degree of control over that data and its use”. In this article, we refer to this latter definition of privacy.
Trustworthy AI is built on the foundational principle of explainability, which supports the gaining of insights into the decision making processes of black-box AI systems (Tabassi, 2023). However, the relationship between privacy and explainability has contrasting aspects. On the one hand, explainability aids privacy in several ways such as in creating privacy awareness in users (Brunotte et al., 2021), in ascertaining that privacy of a system is achieved (Doshi-Velez & Kim, 2017; Müftüoğlu et al., 2022), and in determining correlations with identifiable data for removal (Hohman et al., 2019). On the other hand, explanations can reveal sensitive information contained in models and training data (Harder et al., 2020; Rawal et al., 2022; Zhao et al., 2021) thus leading to privacy risks (Kuppa & Le-Khac, 2021). Thus, there are conflicting outcomes (Guerra-Manzanares et al., 2023; Nguyen et al., 2023; Sanderson et al., 2023; Spartalis et al., 2023) of including explainability as a non-functional requirement in AI systems.

### v3

@@ 2.3.4 Feature-based methods
These explanation methods score or measure the effect of individual input features on the output of the model (arrietaExplainableArtificialIntelligence2020; bhattEvaluatingAggregatingFeaturebased2020; dwivediExplainableAIXAI2023; strobelDataPrivacyTrustworthy2022). They are also referred to as feature importance (mcdermidArtificialIntelligenceExplainability2021), feature relevance (arrietaExplainableArtificialIntelligence2020) or attribution-based (liuPleaseTellMe2024) methods. They are based on the attribution problem which is the distribution of the output of a model for a specific input to its base features (sundararajanManyShapleyValues2020). Two important categories of feature-based methods identified in literature are perturbation and backpropagation-based methods (anconaBetterUnderstandingGradientbased2018; mcdermidArtificialIntelligenceExplainability2021).
  • Perturbation-based methods remove, alter, or mask an input feature or set of features and observe the difference with the original output (mcdermidArtificialIntelligenceExplainability2021). Some perturbation-based methods are LIME (ribeiroWhyShouldTrust2016), permutation feature importance (breimanRandomForests2001), SHAP (lundbergUnifiedApproachInterpreting2017) and MASK (fongInterpretableExplanationsBlack2017).
  • Backpropagation-based methods compute input attributions in forward and backward passes of the network (anconaBetterUnderstandingGradientbased2018). The use of the gradient of the output with the respective input features (mcdermidArtificialIntelligenceExplainability2021; strobelDataPrivacyTrustworthy2022) is a common approach in these methods and is referred to as gradient-based approach. Methods used on images that determine the global importance of pixels, generate saliency maps, and are referred to as pixel-level attribution methods (kapishnikovXRAIBetterAttributions2019; molnarInterpretableMachineLearning2023). Some examples of backpropagation-based methods are gradient (simonyanDeepConvolutionalNetworks2014), gradient x input (shrikumarNotJustBlack2017), guided backpropagation (springenbergStrivingSimplicityAll2015) and integrated gradients (sundararajanAxiomaticAttributionDeep2017).
@@ 2.3.5 Concept-based methods
Concept-based methods aim to uncover high-level concepts above features, pixels and characters (ghorbani2019towards), that influence the behaviour of a model (bereska2024mechanistic). The use of human understandable concepts result in naturally interpretable explanations. Desiderata such as meaningfulness, coherence and importance are proposed for this category of explanations (ghorbani2019towards). Concept activation vectors (CAVs) (kim2018interpretability) is a concept-based method that leverages linear classifiers to determine the presence or absence of concepts corresponding to a set of input examples. The use of concept activation regions (crabbe2022concept) enhances CAVs and allows the underlying examples to be distributed across the latent space of the model, resulting in the discovery of global explanations. In another approach, concept bottleneck models (koh2020concept) predict intermediate human-specified concepts in the model and utilise these for making predictions. The models also support human interventions on these concepts and facilitate the generation of counterfactuals, thus enabling human-model collaboration (koh2020concept).
@@ 2.3.6 Probing-based methods
Probing-based methods use classifiers or probes to determine the knowledge captured by a model (sajjad2022neuron; schneider2024explainable). The technique involves using the internal activations of networks to train probing classifiers for prediction of embedded properties (alain2017probing; belinkov2022probes). However, such probing experiments require the properties to be known apriori along with clear specifications of the original/probing tasks, the models and datasets used. The outcomes of these experiments can be applied for tuning the model on certain tasks or determining the encoded information for downstream tasks.
@@ 2.3.7 Neuron activation methods
These methods focus on the behaviour of neurons that are responsible for specific outcomes or represent learned linguistic properties (zhaoExplainabilityLargeLanguage2024). Proposed methods in this category, such as linguistic correlation analysis and cross-model correlation analysis, identify and analyse such neurons for post-hoc explainability (dalvi2019neuronact). Individual or groups of neurons may be discovered as salient for a property and ranked based their importance to the model’s task (dalvi2019neuronact). These methods may also experience trade-offs between accuracy and selectivity (antverg2022neuronact).
@@ 2.3.8 Mechanistic interpretability
Mechanistic interpretability (bereska2024mechanistic) seeks to understand the model outputs by reverse engineering (olah2022mechanistic) the system. Analogous to understanding complex computer programs, this approach aims at dissecting network activations into independently understandable units. It considers the analysis of individual components of the system, such as features, connections and neurons, to determine the causes of outputs. This approach to interpretability includes techniques that could be applied before, during or after the training process. Techniques, such as feature visualisation (zimmermann2021mechint) and sparse autoencoders (cunningham2024sae), are some of the methods from this category.
@@
Example-based Anchors Model-agnostic ribeiroAnchorsHighPrecisionModelAgnostic2018
Contrastive explanations Model-agnostic dhurandharExplanationsBasedMissing2018
Counterfactuals Model-agnostic wachterCounterfactualExplanationsOpening2017
Influence functions Model-agnostic kohUnderstandingBlackboxPredictions2017
Prototypes and criticisms Model-agnostic kimExamplesAreNot2016
Knowledge-based Semantic web technologies Model-agnostic seeligerSemanticWebTechnologies2019
Neuro-symbolic approaches Model-specific hitzlerNeurosymbolicApproachesArtificial2022
Feature-based Perturbation-based LIME Model-agnostic ribeiroWhyShouldTrust2016
Permutation Feature Importance Model-agnostic breimanRandomForests2001
SHAP Model-agnostic lundbergUnifiedApproachInterpreting2017
MASK Model-agnostic fongInterpretableExplanationsBlack2017
Backpropagation-based Gradient Model-specific simonyanDeepConvolutionalNetworks2014
Gradient x Input Model-specific shrikumarNotJustBlack2017
Guided Backpropagation Model-specific springenbergStrivingSimplicityAll2015
Integrated Gradients Model-specific sundararajanAxiomaticAttributionDeep2017
Concept-based Concept activation vectors Model-specific kim2018interpretability
Concept bottleneck models Model-specific koh2020concept
Concept activation regions Model-specific crabbe2022concept
Probing-based Probing experiments Model-specific alain2017probing
Neuron activation Linguistic correlation analysis Model-specific dalvi2019neuronact
Cross-model correlation analysis Model-specific dalvi2019neuronact
Mechanistic interpretability Feature visualisation Model-specific zimmermann2021mechint
Sparse autoencoders Model-specific cunningham2024sae
@@ 2.4 Related reviews
An in-depth overview of the core concepts and taxonomies in XAI was provided by arrietaExplainableArtificialIntelligence2020. mohseniMultidisciplinarySurveyFramework2021 conducted an interdisciplinary survey and proposed a comprehensive framework for design and evaluation of XAI methods. dwivediExplainableAIXAI2023 covered a wide breadth of explanation algorithms, programming frameworks and software toolkits for XAI development. aliExplainableArtificialIntelligence2023 examined explainability through the lens of trustworthiness detailing evaluation metrics, available software packages and XAI datasets. bodriaBenchmarkingSurveyExplanation2023 systematically categorized explanation methods and benchmarked prominent methods using quantitative metrics. muralidharElementsThatInfluence2023 reviewed transparency elements from human computer interaction (HCI) in the context of explanations while cambriaSurveyXAINatural2023 investigated presentation methods and usage of natural language with XAI.
@@
Beyond these broad surveys, domain specific reviews have also emerged. For example, XAI in healthcare has been surveyed by payrovnaziriExplainableArtificialIntelligence2020 and yangUnboxBlackboxMedical2022; in cybersecurity by capuanoExplainableArtificialIntelligence2022 and in energy and power systems by machlevExplainableArtificialIntelligence2022. Methodology focussed reviews also exist, covering counterfactuals (guidottiCounterfactualExplanationsHow2022), data-driven knowledge-aware XAI systems (liSurveyDatadrivenKnowledgeaware2020), knowledge-graph based XAI (rajabiKnowledgegraphbasedExplainableAI2022; tiddiKnowledgeGraphsTools2022) and semantic web technologies for explanations (seeligerSemanticWebTechnologies2019). Recent advances include the intersection of explainability and federated learning (FL), termed as Federated XAI (Fed-XAI), reviewed by lopez-blancoFederatedLearningExplainable2023 and categorisation of explanation techniques for transformer-based language models based on training paradigms as surveyed by zhaoExplainabilityLargeLanguage2024. A comprehensive review of mechanistic interpretabililty in LLMs for achieving AI safety was conducted by bereska2024mechanistic.
@@
Recent literature has increasingly highlighted the potential of malicious exploitation of XAI interfaces. baniecki2024adversarial present a survey of adversarial attacks in XAI and discuss corresponding defenses. While their study addresses a specific type of XAI privacy attack, its primary emphasis is on non-privacy attacks including data poisoning, backdoor attacks, model manipulation and attacks on fairness metrics through explanations. A categorisation of XAI-aware attacks was proposed in another recent work (Noppel2024adversarial; Noppel2024SoK) based on the outputs that are preserved in the attack. The work also categorised attacks based on the scope of explanation alteration and adversary capabilities. However, this work targets the robustness of post-hoc XAI and though the proposed categorisation is relevant to privacy attacks, it does not directly focus on the privacy aspect of XAI safety.
@@
The focus of this review diverges from prior reviews by specifically examining the privacy risks that target the identification and exposure of personal or sensitive information through the misuse of XAI interfaces in AI systems. Further, we review strategies used by researchers in mitigating the privacy leakage in XAI. An early work by spartalis2023balancing, provides a short review on XAI privacy risks and applicable defense mechanisms. The review also identifies security risks such as evasion and poisoning through XAI interfaces. Though this work raises awareness for further research in this area, it mainly focuses on early privacy attacks and discusses limited defense mechanisms. A recent review (nguyen2025privacy) explores the intersection of privacy and explainability through attacks and countermeasures, but differs from our review in its methodology and lacks a detailed examination of the core requirements underpinning the design of privacy preserving XAI. Our review employs an established scoping review methodology guided by clearly defined research questions. The resulting taxonomy of XAI privacy risks and corresponding mitigation methods are distilled from the understanding of existing literature across the privacy and XAI communities. This methodology enables a structured and rigorous approach to addressing the research questions through the analysis of the selected studies.
@@ 3 Method
We conducted a scoping review based on the Preferred Reporting Items for Systematic reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) (triccoPRISMAExtensionScoping2018). This section elaborates the process followed, the proposed taxonomy of XAI privacy risks and the identified research trends.
@@ 3.1 Literature selection and extraction
A 4-step process was employed comprising of identification, screening, eligibility, and extraction, as illustrated in Figure 2. In the initial step of identification, Elsevier Engineering Village (elsevierb.v.EngineeringVillage) search platform was used and the search was conducted on Compendex and Inspec databases. These databases index publications from leading computer science publishers, including IEEE, ACM, Springer and Elsevier. A researcher formulated the search string, using the two main concepts of privacy and explainability, with the help of a librarian. This was applied on the title, subject, and abstract fields. The reseacher worked with the librarian to set the search, inclusion and exclusion criteria as detailed below for reproducibility:
@@
  • Period of publication: January 1, 2019, to December 31, 2024. The start year was chosen based on the seminal works (milliModelReconstructionModel2019; shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021) published on this topic.
@@
The search results comprising of 3,766 studies were exported from Engineering Village and imported into Covidence (Covidence) review management software by one researcher. During the import process, the software merged duplicate studies from the databases, retaining only unique records. After deduplication, 1,943 studies were forwarded for screening wherein the title and abstract were examined by the researcher to determine relevance to RQ1 or RQ2 while considering the inclusion criteria. Out of 1,943 studies, 69 studies were moved to the next step to determine eligibility wherein the full text of the identified articles were examined with respect to RQ1 and RQ2. This step was conducted by one researcher and independently verified by a second researcher. Since the focus of this article is privacy leakage through XAI interfaces, studies that addressed adversarial XAI (such as backdoor, poisoning, etc.) without overlap on privacy were eliminated. Studies that discussed general privacy issues in ML without XAI context were also eliminated. Related survey papers that appeared in the search results were not extracted but used to determine any additional studies relevant to the research questions. These steps resulted in removal of 17 studies and addition of 5 studies identified through forward and backward searches. The removal and addition of studies were discussed and verified by two researchers. Thus, overall the 4-step scoping review process resulted in extraction of 57 studies.
@@ 4 Privacy Risks in XAI
Traditionally privacy is referred to as the “right to be left alone” (warrenRightPrivacy1890) and the “claim of individuals, groups, or institutions to determine for themselves when, how, and to what extent information about them is communicated to others” (westinPrivacyFreedom1967). In the modern context, with availability, collection, and collation of copious information about individuals through online and offline sources, the concept of information privacy is more applicable and refers to the ability of individuals to exert control on their own data (curzonPrivacyArtificialIntelligence2021). clarkeInternetPrivacyConcerns1999 has defined information privacy as the “claims of individuals that data about themselves should generally not be available to other individuals and organizations, and that, where data is possessed by another party, the individual must be able to exercise a substantial degree of control over that data and its use”. In this article, we refer to this latter definition of privacy.
Trustworthy AI is built on the foundational principle of explainability, which supports the gaining of insights into the decision making processes of black-box AI systems (tabassiAIRiskManagement2023). However, the relationship between privacy and explainability has contrasting aspects. On the one hand, explainability aids privacy in several ways such as in creating privacy awareness in users (brunotteCanExplanationsSupport2021), in ascertaining that privacy of a system is achieved (doshi-velezRigorousScienceInterpretable2017; muftuogluPrivacyPreservingMechanismsExplainability2022), and in determining correlations with identifiable data for removal (hohmanGamutDesignProbe2019). On the other hand, explanations can reveal sensitive information contained in models and training data (harderInterpretableDifferentiallyPrivate2020; rawalRecentAdvancesTrustworthy2022; zhaoExploitingExplanationsModel2021) thus leading to privacy risks (kuppaAdversarialXAIMethods2021). Thus, there are conflicting outcomes (guerra-manzanaresPrivacyPreservingMachineLearning2023; nguyenXRandDifferentiallyPrivate2023; sandersonImplementingResponsibleAI2023; spartalis2023balancing) of including explainability as a non-functional requirement in AI systems.